OpenStudy (noahbred):

Please help? Which mathematical property is demonstrated? 5 • 7 • 4 = 5 • 4 • 7 A. commutative property of addition B. commutative property of multiplication C. associative property of addition D. associative property of multiplication

OpenStudy (shadowlegendx):

Well, you can ignore the ones having to do with addition, as here, they are multiplying

OpenStudy (shadowlegendx):

The associative property states that you can add or multiply regardless of how the numbers are grouped. By 'grouped' we mean 'how you use parenthesis'. In other words, if you are adding or multiplying it does not matter where you put the parenthesis. Add some parenthesis anywhere you like!. The word "commutative" comes from "commute" or "move around", so the Commutative Property is the one that refers to moving stuff around. For addition, the rule is "a + b = b + a"; in numbers, this means 2 + 3 = 3 + 2. For multiplication, the rule is "ab = ba"; in numbers, this means 2×3 = 3×2.

OpenStudy (shadowlegendx):

So with A and C crossed out, what do you think the answer is?

OpenStudy (noahbred):

D

OpenStudy (shadowlegendx):

Are you sure?

OpenStudy (shadowlegendx):

Why do you think itʻs D?

OpenStudy (noahbred):

because it has multipulcation in the answer and we are not adding so its not the other asnwer

OpenStudy (shadowlegendx):

B. commutative property of multiplication D. associative property of multiplication both have multiplication One of them is the answer

OpenStudy (shadowlegendx):

The associative property is about parentheses, and how they are grouped. However, the commutative property is about how numbers can be moved around. What differences do we see from each side of the equation? 5 • 7 • 4 = 5 • 4 • 7

OpenStudy (noahbred):

they are switched so its b hahahaha thanks

OpenStudy (shadowlegendx):

np

OpenStudy (noahbred):

Which mathematical property is demonstrated? If x = –3 and –3 = z, then x = z. A. transitive property of equality B. closure property of addition C. symmetric property of equality D. closure property of multiplication Which mathematical property is demonstrated? (25 + 84) + 16 = 25 + (84 + 16) A. commutative property of addition B. associative property of addition C. associative property of multiplication D. commutative property of multiplication Which mathematical property is demonstrated? (25 + 84) + 16 = 25 + (84 + 16) A. commutative property of addition B. associative property of addition C. associative property of multiplication D. commutative property of multiplication